# An Ion-Intercalation Memristor for Enabling Full Parallel Writing in Crossbar Networks

### v2

@@ I Introduction
The sneak path problem originates from a fundamental structural issue. By the design of the crossbar network, memristors on the same row share the same voltage rail, which is intended to enable efficient multiply-and-accumulate (MAC) operations during the readout stage. On the other hand, in conventional two-terminal memristor designs, read and write signals share the same physical path. Thus, both readout and write-in stages share the same circuit topology. During the write-in stage, the shared voltage rail inevitably exposes non-target memristors to the same write-in voltage, forming a parallel circuit with multiple unintended current paths around the target cell, the essence of the sneak path problem. Although approaches like 1T1R or 1S1R introduce partial access control to suppress sneak currents, they only provide semi-independent switching because one end of the write path remains shared across rows or columns, inherently limiting the scalability of simultaneous multi-cell programming. Consequently, only one switch can be active per row. Thus, for a network with a size of $M\times M$ cells, updating the entire network requires either iterating through each individual cell (${\cal O}(M^{2})$), or row-wise parallel (${\cal O}(M)$) [22].
@@
We propose a memristor that leads to architecture revolution for in-memory computing using memristor arrays. Unlike conventional filament-based memristors that rely on nanoscale defect migration and suffer from stochastic drift, our device uses a statistical resistivity model that leverages ion intercalation into a polymer buffer layer, where the overall ion concentration governs conductance. The bulk statistical property of this modulation ensures high stability and reproducibility; the structure physically separates the read and write paths to ensure read-write orthogonality, eliminating interference between computational and programming signals and enabling true parallel write capability. Furthermore, it supports a parallel programmable network via digital logic and independent control signals, reducing write time complexity from ${\cal O}(M^{2})$ to ${\cal O}(1)$ regardless of array size.
@@ III Methodology
This concept is physically realized in the device structure shown in Fig. 2, where directional ionic programming and perpendicular electrical readout are implemented using a layered system with a polymeric buffer layer.
@@ III-A Electrical Properties of the Memristor Designed
During the write cycle, lithium ions in the electrolyte are driven by an electric field into the buffer layer (Fig. 3). This intercalation increases the polymer’s conductivity by introducing additional charge carriers, thereby reducing its resistivity. During discharge, the reversed field extracts lithium ions from the buffer, increasing its resistivity. By varying the duration of charging and discharging, the lithium ion concentration (thus the resistance) can be fine-tuned. This enables continuous, reversible resistance programming, allowing the device to function as a memristor with adjustable states.
@@
In Chua’s memristor formulation, the magnetic flux $\phi(t)$ is a function of the total transported charge $q(t)$, and the memristance is defined as its derivative [7]: $M(q)=d\phi/dq$. A programming voltage $V_{p}$ drives lithium ions into the polymer buffer layer, generating a vertical electric field $E=V_{p}/d$, where $d$ is the cathode-to-anode distance. The ions drift velocity is $v=\mu E=\mu V_{p}/d$ ($\mu$ is the ionic mobility). Assuming a uniform initial ion concentration $c_{0}$ in the electrolyte, the resulting ionic current density magnitude is
@@
where $A=l_{x}l_{z}$ is the area of the cross section, $l_{z}$ is the height of the readout electrode, $l_{x}$ is the distance between readout electrodes, and $l_{y}$ is the thickness of the polymer layer.
The total charge injected into the polymer layer over time $t$ is $q(t)=I_{p}t$. These ions accumulate within the buffer volume $Al_{y}$ establishing a local concentration
@@
The polymer-buffer’s conductivity linearly depends on $c(t)$:
@@
where $\mu_{e}$ is the effective electronic mobility within the polymer matrix, representing the average drift mobility of charge carriers generated by ion intercalation. This parameter governs the final conductivity of the doped region under an applied readout field. The corresponding resistance is expressed as
@@ III-B2 Read Operation
At the device level, in the read mode the memristor only experiences an electric field along the read direction. Importantly, this electric field is orthogonal to the programming direction and therefore does not induce ion migration out of the buffer layer. As depicted in Figs. 7 and 3, the orthogonality of read and write current paths ensures that the read process does not disturb the programmed conductance state. This design feature significantly mitigates the risk of long-term conductance drift triggered by repeated read operations.
@@ IV-C Long-term Retention Characterization
Fig. 10 shows the time-dependent resistance relaxation behavior of the device S2 after the programming stage. As illustrated, the resistance gradually increases during the retention window and returns close to its initial value within approximately 48 hours. However, the resistance evolution is fully reversible, since the resistance modulation arises from the intercalation and deintercalation of mobile ions in the polymer buffer layer. The return of the resistance to its initial state demonstrates that the programming process attributed to bulk ionic-doping mechanism, not a generating permanent structure. The current result already validates the reversibility of the underlying physical process. However, further material and interface optimization is needed.
@@ V-C Routing Overhead as an Inevitable Cost of Parallelism
VI Conclusion

### v3

@@ I Introduction
The sneak path problem originates from a fundamental structural issue. By the design of the crossbar network, memristors on the same row share the same voltage rail, which is intended to enable efficient multiply-and-accumulate (MAC) operations during the readout stage. On the other hand, in conventional two-terminal memristor designs, read and write signals share the same physical path. Thus, both readout and write-in stages share the same circuit topology. During the write-in stage, the shared voltage rail inevitably exposes non-target memristors to the same write-in voltage, forming a parallel circuit with multiple unintended current paths around the target cell, the essence of the sneak path problem. Although approaches like 1T1R or 1S1R introduce partial access control
to suppress sneak currents, they only provide semi-independent switching because one end of the write path remains shared across rows or columns, inherently limiting the scalability of simultaneous multi-cell programming. Consequently, only one switch can be active per row. Thus, for a network with a size of $M\times M$ cells, updating the entire network requires either iterating through each individual cell (${\cal O}(M^{2})$), or row-wise parallel (${\cal O}(M)$) [22].
@@
We propose a memristor that leads to architecture revolution for in-memory computing using memristor arrays. Unlike conventional filament-based memristors that rely on nanoscale defect migration and suffer from stochastic drift, our device uses a statistical resistivity model that leverages ion intercalation into an ion acceptor layer, where the overall ion concentration governs conductance. The bulk statistical property of this modulation ensures high stability and reproducibility; the structure physically separates the read and write paths to ensure read-write orthogonality, eliminating interference between computational and programming signals and enabling true parallel write capability. Furthermore, it supports a parallel programmable network via digital logic and independent control signals, reducing write time complexity from ${\cal O}(M^{2})$ to ${\cal O}(1)$ regardless of array size.
@@ III Methodology
This concept is physically realized in the device structure shown in Fig. 2, where directional ionic programming and perpendicular electrical readout are implemented using a layered system with an ion acceptor layer.
@@ III-A Electrical Properties of the Memristor Designed
During the write cycle, lithium ions in the electrolyte are driven by an electric field into the ion acceptor layer (Fig. 3). This intercalation increases the ion acceptor layer’s conductivity by introducing additional charge carriers, thereby reducing its resistivity. During discharge, the reversed field extracts lithium ions from the ion acceptor layer, increasing its resistivity. By varying the duration of charging and discharging, the lithium ion concentration (thus the resistance) can be fine-tuned. This enables continuous, reversible resistance programming, allowing the device to function as a memristor with adjustable states.
@@
In Chua’s memristor formulation, the magnetic flux $\phi(t)$ is a function of the total transported charge $q(t)$, and the memristance is defined as its derivative [7]: $M(q)=d\phi/dq$. A programming voltage $V_{p}$ drives lithium ions into the ion acceptor layer, generating a vertical electric field $E=V_{p}/d$, where $d$ is the cathode-to-anode distance. The ions drift velocity is $v=\mu E=\mu V_{p}/d$ ($\mu$ is the ionic mobility). Assuming a uniform initial ion concentration $c_{0}$ in the electrolyte, the resulting ionic current density magnitude is
@@
where $A=l_{x}l_{z}$ is the area of the cross section, $l_{z}$ is the height of the readout electrode, $l_{x}$ is the distance between readout electrodes, and $l_{y}$ is the thickness of the ion acceptor layer.
The total charge injected into the ion acceptor layer over time $t$ is $q(t)=I_{p}t$. These ions accumulate within the ion acceptor volume $Al_{y}$ establishing a local concentration
@@
The ion acceptor layer’s conductivity linearly depends on $c(t)$:
@@
where $\mu_{e}$ is the effective electronic mobility within the ion-acceptor matrix, representing the average drift mobility of charge carriers generated by ion intercalation. This parameter governs the final conductivity of the doped region under an applied readout field. The corresponding resistance is expressed as
@@ III-B2 Read Operation
At the device level, in the read mode the memristor only experiences an electric field along the read direction. Importantly, this electric field is orthogonal to the programming direction and therefore does not induce ion migration out of the ion acceptor layer. As depicted in Figs. 7 and 3, the orthogonality of read and write current paths ensures that the read process does not disturb the programmed conductance state. This design feature significantly mitigates the risk of long-term conductance drift triggered by repeated read operations.
@@ IV-C Long-term Retention Characterization
Fig. 10 shows the time-dependent resistance relaxation of device S2 after programming. Under zero-bias conditions, the resistance gradually increases and returns close to its initial value within approximately 48 hours.
This relaxation behavior is primarily determined by the material system used in the present prototype. The ion acceptor layer is based on polyethylene glycol (PEG), where lithium ions are stored through relatively weak coordination interactions. As a result, the ionic potential wells are shallow, and thermal activation enables slow back-diffusion of ions once the external programming field is removed. This leads to a gradual reduction in carrier concentration and a corresponding resistance increase over long time scales.
The resistance evolution is fully reversible, and the device remains reprogrammable after relaxation. This behavior is consistent with a bulk ionic redistribution mechanism rather than irreversible filament formation.
@@ V-A On the Necessity of Array-Level Sneak-Path Validation
We note that in practical implementations, non-idealities (e.g., parasitics, unintended coupling, or fabrication defects) may cause the realized topology to deviate from the ideal circuit assumption and can introduce additional failure modes. A systematic array-level characterization of such non-ideal effects is beyond the scope of this paper .
@@ V-C Routing Overhead as an Inevitable Cost of Parallelism
V-D Impact of Device Geometry on Programming Speed
The relatively slow programming speed observed in the present prototype should be interpreted in the context of its intentionally oversized physical dimensions. The proof-of-concept device reported in this work has a characteristic programming length scale on the order of 2 cm, which is several orders of magnitude larger than that of state-of-the-art nanoscale memristors.
In the proposed device, resistance modulation is governed by a field-driven ion intercalation process. Under this mechanism, the characteristic programming time is dominated by ionic drift rather than electronic transport. The ion drift velocity is given by $v=\mu Vp/d$, where $\mu$ is the ionic mobility. Vp is the programming voltage, and d is the effective ion transport distance. As a result, the time required for ions to traverse the active region scales linearly with d.
For nanoscale filamentary or interfacial memristors, d typically ranges from a few nanometers to tens of nanometers. In contrast, the present prototype deliberately employs a macroscopic transport distance on the order of centimeters to simplify fabrication, enable direct probing, and validate the proposed physical principle. This difference in length scale alone introduces a slowdown exceeding ten orders of magnitude when compared to nanoscale devices, even under comparable electric fields.
Importantly, this reduced programming speed is not an intrinsic limitation of the ion-intercalation-based mechanism itself. Rather, it is a direct and predictable consequence of the macroscopic geometry chosen for this proof-of-concept implementation. The analytical model developed in Section III-A captures this dependence through the geometric parameters, and the experimental results are consistent with a drift-dominated transport regime.
When the same device structure is scaled down to the nanometer regime through materials engineering, the programming time can be substantially reduced. By shrinking the ion transport distance via nanoscale integration of the reservoir, electrolyte, and acceptor layers, the characteristic drift length d is reduced by several orders of magnitude, directly accelerating the ion-intercalation process under the same electric field. Importantly, such scaling does not require any change to the underlying operating principle of the device, but rather relies on established thin-film deposition, solid-state electrolyte engineering, and interface optimization techniques. Therefore, the present prototype should be understood as a macroscopic validation of the physical mechanism and architectural concept, while practical high-speed operation is expected to emerge naturally as the structure is scaled to nanometer dimensions through material and process optimization.
VI Future Work
While this work focuses on validating the physical mechanism and architectural feasibility of a memristor with orthogonal read and write paths, several directions remain for further development toward practical deployment.
First, future efforts will focus on identifying fully solid-state material systems through materials engineering. Replacing liquid or gel-based components with solid electrolytes and solid ion reservoirs is expected to improve device robustness, compatibility with microfabrication processes, and long-term stability. Solid-state implementations would also enable tighter geometric control of ion transport paths, which is essential for large-scale integration.
Second, retention performance can be further improved by engineering deeper ionic potential wells within the acceptor layer. By selecting materials and interfaces with stronger ion binding energies or higher activation barriers for back-diffusion, spontaneous relaxation of the programmed state can be suppressed. Such approaches are well established in solid-state electrochemistry and are directly applicable to the proposed device structure without altering its operating principle.
Third, device miniaturization to the nanometer scale is expected to significantly reduce programming time. In the present prototype, programming dynamics are governed by drift-dominated ion transport across a macroscopic distance. Scaling the structure down to nanometer dimensions through material and process optimization will proportionally reduce the ion transport length. Based on linear extrapolation from the drift model and the design parameters of the MEMS-compatible implementation, programming latency is expected to decrease by multiple orders of magnitude while maintaining the same voltage range and control methodology.
Finally, large-scale fabrication and statistical characterization will be necessary to evaluate device-to-device variability and programming consistency. Batch fabrication using MEMS or CMOS-compatible processes will enable systematic assessment of uniformity, yield, and endurance across large arrays. Such studies are critical for validating the scalability claims of the proposed architecture and for establishing confidence in its applicability to parallel in-memory computing systems.
VII Conclusion